# Beyond the Hodge Theorem: curl and asymmetric pseudodifferential projections

### v3

@@ Appendix A Exterior calculus
We define the action of the Hodge star $*:\Omega^{k}(M)\to\Omega^{3-k}(M)$ on a rank $k$ antisymmetric tensor as
@@
Appendix B The spectrum of the Laplacian on a Berger sphere
@@
In preparation for the calculations that will appear in Appendix C, we list here, in the form of a theorem, the eigenvalues of the Laplacian on a Berger sphere. These were first obtained in [52, Lemma 4.1] and reproduced in [41, Theorem 5.5] and [38, Proposition 3.9].
B.1 Definitions and notation
@@
The above definition of orientation agrees with that from [27, Appendix A]. Spherical coordinates
@@
Of course, if one is only interested in the study of the Laplace–Beltrami operator, orientation is irrelevant. However, with $\operatorname{curl}$ or the Dirac operator [16, subsection 9.1] prescribing orientation of the manifold is essential.
@@
$\mathbf{V}_{1}:=-\textbf{x}^{4}\dfrac{\partial}{\partial\textbf{x}^{1}}-\textbf{x}^{3}\frac{\partial}{\partial\textbf{x}^{2}}+\textbf{x}^{2}\dfrac{\partial}{\partial\textbf{x}^{3}}+\textbf{x}^{1}\dfrac{\partial}{\partial\textbf{x}^{4}}\,,$ (B.1)
$\mathbf{V}_{2}:=\phantom{-}\textbf{x}^{3}\dfrac{\partial}{\partial\textbf{x}^{1}}-\textbf{x}^{4}\frac{\partial}{\partial\textbf{x}^{2}}-\textbf{x}^{1}\dfrac{\partial}{\partial\textbf{x}^{3}}+\textbf{x}^{2}\dfrac{\partial}{\partial\textbf{x}^{4}}\,,$
$\mathbf{V}_{3}:=-\textbf{x}^{2}\dfrac{\partial}{\partial\textbf{x}^{1}}+\textbf{x}^{1}\frac{\partial}{\partial\textbf{x}^{2}}-\textbf{x}^{4}\dfrac{\partial}{\partial\textbf{x}^{3}}+\textbf{x}^{3}\dfrac{\partial}{\partial\textbf{x}^{4}}\,.$
[…]
@@
$B_{\mathrm{prin}}(y,\eta)=-g^{\alpha\beta}(y)\eta_{\alpha}\eta_{\beta}$ (B.3)
@@
$\begin{split}V_{1}&=\sqrt{1-|y|^{2}}\dfrac{\partial}{\partial y^{1}}-y^{3}\dfrac{\partial}{\partial y^{2}}+y^{2}\dfrac{\partial}{\partial y^{3}},\\ V_{2}&=y^{3}\dfrac{\partial}{\partial y^{1}}+\sqrt{1-|y|^{2}}\dfrac{\partial}{\partial y^{2}}-y^{1}\dfrac{\partial}{\partial y^{3}},\\ V_{3}&=-y^{2}\dfrac{\partial}{\partial y^{1}}+y^{1}\dfrac{\partial}{\partial y^{2}}+\sqrt{1-|y|^{2}}\dfrac{\partial}{\partial y^{3}}.\end{split}$ (B.4)
Formulae (B.2), (B.3) and (B.4) imply
$\footnotesize g^{\alpha\beta}(y)=\begin{pmatrix}1-(y^{1})^{2}-\frac{a^{2}-1}{a^{2}}(y^{2})^{2}&-\frac{y^{1}y^{2}}{a^{2}}&-y^{1}y^{3}+\frac{a^{2}-1}{a^{2}}y^{2}\sqrt{1-|y|^{2}}\\ -\frac{y^{1}y^{2}}{a^{2}}&1-(y^{2})^{2}-\frac{a^{2}-1}{a^{2}}(y^{1})^{2}&-y^{2}y^{3}-\frac{a^{2}-1}{a^{2}}y^{1}\sqrt{1-|y|^{2}}\\ -y^{1}y^{3}+\frac{a^{2}-1}{a^{2}}y^{2}\sqrt{1-|y|^{2}}&-y^{2}y^{3}-\frac{a^{2}-1}{a^{2}}y^{1}\sqrt{1-|y|^{2}}&\frac{a^{2}-1}{a^{2}}(y^{2})^{2}+\frac{a^{2}-1}{a^{2}}(y^{2})^{2}-\frac{1-(y^{3})^{2}}{a^{2}}\\ \end{pmatrix}.$ (B.5)
@@
    $\widetilde{V}_{1}:=V_{1}\,,\qquad\widetilde{V}_{2}:=V_{2}\,,\qquad\widetilde{V}_{3}:=\frac{1}{a}V_{3}$ (B.6)
    are orthonormal with respect to the metric $g$ defined by (B.5).
@@
    The operator $B$ defined in accordance with (B.2) is the Laplace–Beltrami operator (henceforth denoted by $\Delta$) on the Berger sphere.
@@
    $\operatorname{curl}_{s}=\begin{pmatrix}\frac{2}{a}&-\frac{1}{a}V_{3}&V_{2}\\ \frac{1}{a}V_{3}&\frac{2}{a}&-V_{1}\\ -V_{2}&V_{1}&2a\\ \end{pmatrix}.$ (B.7)
    Here the subscript $\,s\,$ indicates that the operator appearing in (B.7) acts on 3-columns of scalar functions. The latter can be obtained by projecting 1-forms onto the orthonormal framing (B.6). The operators $\operatorname{curl}$ and $\operatorname{curl}_{s}$ are related as
@@
    See also the projection operator $\mathbf{S}$ in [15, Section 5.1].
B.2 The spectrum
Theorem B.2.
The eigenvalues of the operator $-\Delta$ on the Berger 3-sphere are
$n(n+2)+\left(a^{-2}-1\right)\left(n-2l\right)^{2}\,,\quad n=0,1,2,3,\dots,\quad 0\leq l\leq\left\lfloor\frac{n}{2}\right\rfloor$ (B.8)

### v4

@@ Appendix A Exterior calculus
We define the action of the Hodge star $\Omega^{k}(M)\to\Omega^{3-k}(M)$ on a rank $k$ antisymmetric tensor as
@@
Appendix B Basic properties of the operator curl
In this appendix we provide, in a self-contained fashion, proofs of the fundamental properties of $\operatorname{curl}$ stated in Section 2.
In order to prove the Theorem 2.1, let us introduce an auxiliary elliptic operator: extended curl.
B.1 Extended curl
Definition B.1.
We define extended curl to be the operator
$\operatorname{curl}_{E}:=\begin{pmatrix}\operatorname{curl}&\mathrm{d}\\ \delta&0\end{pmatrix}:(\Omega^{1}\cap H^{1})\oplus(\Omega^{0}\cap H^{1})\to\Omega^{1}\oplus\Omega^{0}\,.$ (B.1)
The operator (B.1) appeared in [52], see also [2, p. 229], [3, pp. 44–45 and p. 66], [4, p. 405], [60, pp. 190–191] and [24, §4].
Remark B.2.
Note that $\operatorname{curl}_{E}$ is an operator of Dirac type, namely, its square is $\delta\mathrm{d}+\mathrm{d}\delta$ on $\Omega^{1}\oplus\Omega^{0}$ (minus the nonpostive Hodge Laplacian). Furthermore, (B.1) is the restriction of the signature operator $\mathrm{d}+\delta$ to differential forms of even degree, in the sense that
$(\operatorname{curl}_{E})_{\mathrm{prin}}(x,\xi)=\begin{pmatrix}[(\operatorname{curl})_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)&i\xi_{\alpha}\\ -ig^{\beta\gamma}(x)\,\xi_{\gamma}&0\end{pmatrix}$. (B.2)
The eigenvalues of (B.2) are
$\pm\|\xi\|\,$, (B.3)
each with multiplicity two, so that we have
$\det\left[(\operatorname{curl}_{E})_{\mathrm{prin}}\right]=\|\xi\|^{4}\,$, (B.4)
Compare formulae (B.3) and (B.4) with (2.3) and (2.2) respectively.
Formulae (B.4) and (B.1) imply that $\operatorname{curl}_{E}$ is elliptic and self-adjoint.
Definition B.3 ([57, Definition 1.7]).
$L(\mathcal{D}(L)\cap V)\subseteq V.$
Proposition B.4.
$\Omega^{1}\oplus\Omega^{0}=[\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1}]\oplus[\delta\Omega^{2}\oplus 0]\oplus[\mathcal{H}^{1}\oplus\mathcal{H}^{0}]$ (B.5)
which are invariant subspaces of the operator $\operatorname{curl}_{E}$ in the sense of Definition B.3. Accordingly, the operator $\operatorname{curl}_{E}$ decomposes into a direct sum of three operators
$\operatorname{curl}_{E}=\operatorname{curl}_{E,\mathrm{d}}\oplus\operatorname{curl}_{E,\delta}\oplus\operatorname{curl}_{E,\mathcal{H}}\,$, (B.6)
where
$\operatorname{curl}_{E,\mathrm{d}}:(\mathrm{d}\Omega^{0}\cap H^{1})\oplus(\delta\Omega^{1}\cap H^{1})\to\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1},$ (B.7)
$\operatorname{curl}_{E,\delta}:(\delta\Omega^{2}\cap H^{1})\oplus 0\to\delta\Omega^{2}\oplus 0,$ (B.8)
$\operatorname{curl}_{E,\mathcal{H}}:\mathcal{H}^{1}\oplus\mathcal{H}^{0}\to\mathcal{H}^{1}\oplus\mathcal{H}^{0}.$ (B.9)
Proof.
That the three subspaces in the RHS of (B.5) are mutually orthogonal and their direct sum gives the whole space $\Omega^{1}\oplus\Omega^{0}$ follows immediately from the Hodge decomposition theorem. Let us show that they are invariant subspaces of $\operatorname{curl}_{E}\,$. We will do this in several steps.
Step 2. The facts that
  • $\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1}]\oplus[\delta\Omega^{2}\oplus 0]$ is the orthogonal complement of $\mathcal{H}^{1}\oplus\mathcal{H}^{0}$
imply that $\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1}]\oplus[\delta\Omega^{2}\oplus 0]$ is an invariant subspace of $\operatorname{curl}_{E}\,$.
$\lambda\in\mathbb{R}.$ (B.10)
$\lambda\neq 0.$ (B.11)
$0=\mu_{0}<\mu_{1}\leq\mu_{2}\leq\dots$ (B.12)
be the eigenvalues of $\Delta\,$ enumerated in increasing order with account of multiplicities. It will be convenient for us to further restrict the choice of $\lambda$ by imposing the condition
$\lambda\not\in\{\pm\sqrt{\mu_{1}}\,,\,\pm\sqrt{\mu_{2}}\,$$\dots\}$. (B.13)
Step 5. In view of (B.10) the operator $\mathcal{R}(\lambda)$ is self-adjoint. The facts that
  • $\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1}]\oplus[\delta\Omega^{2}\oplus 0]$ is an invariant subspace of $\mathcal{R}(\lambda)$ and
$\mathrm{d}\Omega^{k-1}_{\infty}:=\{\mathrm{d}\omega\ |\ \omega\in\Omega^{k-1}_{\infty}\},$ (B.14)
$\delta\Omega^{k+1}_{\infty}:=\{\delta\omega\ |\ \omega\in\Omega^{k+1}_{\infty}\}.$ (B.15)
The Hilbert spaces $\Omega^{k}$, $\mathrm{d}\Omega^{k-1}$ and $\delta\Omega^{k+1}$ are defined as $L^{2}$ closures of the vector spaces $\Omega^{k}_{\infty}$, $\mathrm{d}\Omega^{k-1}_{\infty}$ and $\delta\Omega^{k+1}_{\infty}$ respectively, see [42, Corollary 3.4.2]. This implies that the vector space $\delta\Omega^{2}_{\infty}\oplus 0$ is dense in the Hilbert spaces $\delta\Omega^{2}\oplus 0$. As $\mathcal{R}(\lambda)$ is a bounded operator, in order to prove that $\delta\Omega^{2}\oplus 0$ is an invariant subspace of $\mathcal{R}(\lambda)$ it is sufficient to prove that
$\mathcal{R}(\lambda)](\delta\Omega^{2}_{\infty}\oplus 0)\subseteq\delta\Omega^{2}_{\infty}\oplus 0\,.$ (B.16)
$v:=\mathcal{R}(\lambda)z=\begin{pmatrix}v_{1}\\ v_{0}\end{pmatrix}\in[\mathrm{d}\Omega^{0}_{\infty}\oplus\delta\Omega^{1}_{\infty}]\oplus[\delta\Omega^{2}_{\infty}\oplus 0].$ (B.17)
Here $z_{1}$, $v_{1}$ are 1-forms and $v_{0}$ is a 0-form (scalar). The $v_{1}$ and $v_{0}$ are infinitely smooth because the operator $\operatorname{curl}_{E}$ is elliptic. Formula (B.17) can be rewritten more explicitly as
$\operatorname{curl}v_{1}+\mathrm{d}v_{0}-\lambda v_{1}$ $=z_{1}\,$, (B.18)
$\delta v_{1}-\lambda v_{0}$ $=0\,.$ (B.19)
We have [63, Section 6.8, p. 223]
$\Omega^{k}_{\infty}=\mathrm{d}\Omega^{k-1}_{\infty}\oplus\delta\Omega^{k+1}_{\infty}\oplus\mathcal{H}^{k},$ (B.20)
compare with (1.1). We already know (see Step 2) that the 1-form $v_{1}$ does not contain a contribution from harmonic 1-forms, so formulae (B.20), (B.14) and (B.15) give us
$v_{1}=v_{1\mathrm{d}}+v_{1\delta}\,$, (B.21)
where
$v_{1\mathrm{d}}=\mathrm{d}w_{0}\,$, (B.22)
$v_{1\delta}=\delta w_{2}$ (B.23)
Step 8. Applying the operator $\delta$ to (B.18) and using formulae (B.21)–(B.23) as well as the fact that $\delta z_{1}=0$, we get $\Delta(\lambda w_{0}-v_{0})=0$, which implies
$v_{0}=\lambda w_{0}+C,$ (B.24)
where $C\in\mathbb{R}$ is a constant. Substituting (B.21) and (B.24) into (B.19) and using formulae (B.22), (B.23), we get
$\Delta w_{0}-\lambda^{2}w_{0}=\lambda C.$ (B.25)
In view of (B.11) and (B.13) equation (B.25) has the unique solution $w_{0}=-\lambda^{-1}C$, which immediately implies
$v_{0}=0,\qquad v_{1\mathrm{d}}=0.$ (B.26)
It only remains to observe that formulae (B.17), (B.21), (B.23) and (B.26) imply $v\in\delta\Omega^{2}_{\infty}\oplus 0$, thus proving the inclusion (B.16). ∎
B.2 Decomposition of the spectrum of extended curl
Recall that extended curl (B.1) is a self-adjoint elliptic operator, hence its spectrum is discrete and eigenfunctions infinitely smooth.
Proposition B.5.
The spectrum of $\operatorname{curl}_{E}$ decomposes into three parts as
$\sigma(\operatorname{curl}_{E})=\sigma(\operatorname{curl})\cup\sigma(\operatorname{curl}_{E,\mathrm{d}})\cup\sigma(\operatorname{curl}_{E,\mathcal{H}}),$ (B.27)
where $\operatorname{curl}$, $\operatorname{curl}_{E,\mathrm{d}}$ and $\operatorname{curl}_{E,\mathcal{H}}$ are the operators (2.4), (B.7) and (B.9) respectively. Formula (B.27) is understood as the (disjoint) union of spectra with account of multiplicities.
Proof.
Proposition B.5 is an immediate consequence of Proposition B.4 and the observation that the operator $\operatorname{curl}_{E,\delta}\,$ defined by formulae (B.1) and (B.8) is precisely our original operator $\operatorname{curl}$ defined by formula (2.4). ∎
Proposition B.5 tells us that the spectrum of curl sits inside the spectrum of extended curl and the two spectra differ by $\sigma(\operatorname{curl}_{E,\mathrm{d}})\cup\sigma(\operatorname{curl}_{E,\mathcal{H}})$. So the task at hand is working out the spectra of $\operatorname{curl}_{E,\mathrm{d}}$ and $\operatorname{curl}_{E,\mathcal{H}}\,$. The following two lemmata address this issue.
Lemma B.6.
Proof.
Lemma B.6 is an immediate consequence of formulae (B.1) and (B.9). ∎
Lemma B.7.
We have
$\sigma(\operatorname{curl}_{E,\mathrm{d}})=\{\pm\sqrt{\mu_{1}}\,,\,\pm\sqrt{\mu_{2}}\,$$\dots\},$ (B.28)
where the $\mu_{j}$ are the eigenvalues (B.12) of $\Delta\,$.
Proof.
$\mathrm{d}v_{0}$ $=\lambda\mathrm{d}w_{0}\,$, (B.29)
$-\Delta w_{0}$ $=\lambda v_{0}\,.$ (B.30)
Elementary analysis of (B.29) (B.30) yields (B.28). ∎
Proposition B.5 and Lemmata B.6B.7 tell us that the spectrum of $\operatorname{curl}$ can be recovered from the spectrum of $\operatorname{curl}_{E}\,$, provided one knows the spectrum of the Laplace–Beltrami operator $\Delta$. Moreover, the eta functions of $\operatorname{curl}$ and $\operatorname{curl}_{E}\,$ are the same, because contributions to $\sigma(\operatorname{curl}_{E})$ coming from the Laplace–Beltrami operator are symmetric about zero.
B.3 Proof of Theorem 2.1
Proof of Theorem 2.1.
(a) Standard elliptic theory tells us that the operator $\operatorname{curl}_{E}$ defined by formula (B.1) is self-adjoint in the operator theoretic sense. Proposition B.4 and formulae (B.1) and (B.8) tell us that our original operator $\operatorname{curl}$ defined by formula (2.4) is part of the operator $\operatorname{curl}_{E}\,$; here the concept of ‘part of an operator’ is understood in accordance with [57, Definition 1.7]. Self-adjointness of $\operatorname{curl}$ now follows from [57, Proposition 3.11].
(b) Discreteness of the spectrum of $\operatorname{curl}$ follows from Proposition B.5. As to the statement that the spectrum of $\operatorname{curl}$ is unbounded both from above and from below, we prove it by writing down spectral asymptotics.
Let $\lambda_{k}$, $k\in\mathbb{Z}\setminus\{0\}$, be the eigenvalues of $\operatorname{curl}$. The choice of particular enumeration is irrelevant for our purposes, but what is important is that eigenvalues are enumerated with account of their multiplicity. We define the positive (++) and negative (-) counting functions for $\operatorname{curl}$ as
$N^{\pm}_{\operatorname{curl}}(\lambda):=\begin{cases}0\quad&\text{for}\quad\lambda\leq 0,\\ \sum_{0<\pm\lambda_{k}<\lambda}\,1\quad&\text{for}\quad\lambda>0.\end{cases}$ (B.31)
We define, in a similar fashion, counting functions $N^{\pm}_{\operatorname{curl}_{E}}(\lambda)$ and $N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)$ for the operators $\operatorname{curl}_{E}$ and $\operatorname{curl}_{E,d}$ respectively. Proposition B.5 and Lemma B.6 tell us that
$N^{\pm}_{\operatorname{curl}}(\lambda)=N^{\pm}_{\operatorname{curl}_{E}}(\lambda)-N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda).$ (B.32)
Lemma B.7 tells us that the evaluation of $N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)$ reduces to the evaluation of eigenvalues of the Laplace–Beltrami operator. The latter is a subject which has been extensively studied over the last 100 years. Application of [38, Theorem 1.1] gives the following asymptotic formula for $N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)$ with sharp remainder term estimate:
$N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)=\frac{\operatorname{Vol}(M)}{6\pi^{2}}\lambda^{3}+O(\lambda^{2})\qquad\text{as}\qquad\lambda\to+\infty,$ (B.33)
where $\operatorname{Vol}(M)$ is the Riemannian volume of the manifold.
Application of [39, Theorem 0.1] gives the following asymptotic formula for $N^{\pm}_{\operatorname{curl}_{E}}(\lambda)$:
$N^{\pm}_{\operatorname{curl}_{E}}(\lambda)=\frac{\operatorname{Vol}(M)}{3\pi^{2}}\lambda^{3}+O(\lambda^{2})=2N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)+O(\lambda^{2})\qquad\text{as}\qquad\lambda\to+\infty.$ (B.34)
The factor 2 appears in the RHS of (B.34) because the eigenvalues of the principal symbol of the operator $\operatorname{curl}_{E}\,$, given by formula (B.3), have multiplicity two.
Formulae (B.32)–(B.34) imply Bär’s [7, Theorem 3.6] asymptotic formula
$N^{\pm}_{\operatorname{curl}}(\lambda)=\frac{\operatorname{Vol}(M)}{6\pi^{2}}\lambda^{3}+O(\lambda^{2})\qquad\text{as}\qquad\lambda\to+\infty.$ (B.35)
Formula (B.35) shows that the spectrum of $\operatorname{curl}$ accumulates to $\infty$ and -$\infty$.
(c) Suppose that zero is an eigenvalue of $\operatorname{curl}$. Let
$u\in\delta\Omega^{2}_{\infty}$ (B.36)
$\mathrm{d}u=0.$ (B.37)
But (B.36) implies
$\delta u=0.$ (B.38)
Formulae (B.37) and (B.38) tell us that the 1-form $u$ is harmonic. As $\delta\Omega^{2}\cap\mathcal{H}^{1}=\{0\}$, we conclude that $u=0$, which is a contradiction.
(d) Proposition B.5, Lemmata B.6 and B.7 and part (c) of Theorem 2.1 tell us that zero is an eigenvalue of $\operatorname{curl}_{E}$ of multiplicity $\operatorname{dim}\mathcal{H}^{1}+1$. Let $P_{\mathcal{H}^{1}}$ and $P_{\mathcal{H}^{0}}$ be orthogonal projections onto the spaces of harmonic 1-forms and 0-forms respectively. Put
Appendix C The spectrum of the Laplacian on a Berger sphere
@@
In preparation for the calculations that will appear in Appendix D, we list here, in the form of a theorem, the eigenvalues of the Laplacian on a Berger sphere. These were first obtained in [62, Lemma 4.1] and reproduced in [50, Theorem 5.5] and [45, Proposition 3.9].
C.1 Definitions and notation
@@
The above definition of orientation agrees with that from [31, Appendix A]. Spherical coordinates
@@
Of course, if one is only interested in the study of the Laplace–Beltrami operator, orientation is irrelevant. However, with $\operatorname{curl}$ or the Dirac operator [17, subsection 9.1] prescribing orientation of the manifold is essential.
@@
$\mathbf{V}_{1}=-\textbf{x}^{4}\dfrac{\partial}{\partial\textbf{x}^{1}}-\textbf{x}^{3}\frac{\partial}{\partial\textbf{x}^{2}}+\textbf{x}^{2}\dfrac{\partial}{\partial\textbf{x}^{3}}+\textbf{x}^{1}\dfrac{\partial}{\partial\textbf{x}^{4}}\,$, (C.1)
$\mathbf{V}_{2}=\phantom{-}\textbf{x}^{3}\dfrac{\partial}{\partial\textbf{x}^{1}}-\textbf{x}^{4}\frac{\partial}{\partial\textbf{x}^{2}}-\textbf{x}^{1}\dfrac{\partial}{\partial\textbf{x}^{3}}+\textbf{x}^{2}\dfrac{\partial}{\partial\textbf{x}^{4}}\,$,
$\mathbf{V}_{3}=-\textbf{x}^{2}\dfrac{\partial}{\partial\textbf{x}^{1}}+\textbf{x}^{1}\frac{\partial}{\partial\textbf{x}^{2}}-\textbf{x}^{4}\dfrac{\partial}{\partial\textbf{x}^{3}}+\textbf{x}^{3}\dfrac{\partial}{\partial\textbf{x}^{4}}\,.$
The vector fields (C.1) are tangent to $\mathbb{S}^{3}$, i.e., they are orthogonal to the gradient of $\|\textbf{x}\|$, or, equivalently, $\mathbf{V}_{j}(\|\textbf{x}\|^{2})=0$, $j=1,2,3$. For $j=1,2,3$ let $V_{j}$ be the restriction of $\mathbf{V}_{j}$ to $\mathbb{S}^{3}$.
@@
$B_{\mathrm{prin}}(y,\eta)=-g^{\alpha\beta}(y)\eta_{\alpha}\eta_{\beta}$ (C.3)
@@
$\begin{split}V_{1}&=\sqrt{1-|y|^{2}}\dfrac{\partial}{\partial y^{1}}-y^{3}\dfrac{\partial}{\partial y^{2}}+y^{2}\dfrac{\partial}{\partial y^{3}},\\ V_{2}&=y^{3}\dfrac{\partial}{\partial y^{1}}+\sqrt{1-|y|^{2}}\dfrac{\partial}{\partial y^{2}}-y^{1}\dfrac{\partial}{\partial y^{3}},\\ V_{3}&=-y^{2}\dfrac{\partial}{\partial y^{1}}+y^{1}\dfrac{\partial}{\partial y^{2}}+\sqrt{1-|y|^{2}}\dfrac{\partial}{\partial y^{3}}.\end{split}$ (C.4)
Formulae (C.2), (C.3) and (C.4) imply
$\footnotesize g^{\alpha\beta}(y)=\begin{pmatrix}1-(y^{1})^{2}-\frac{a^{2}-1}{a^{2}}(y^{2})^{2}&-\frac{y^{1}y^{2}}{a^{2}}&-y^{1}y^{3}+\frac{a^{2}-1}{a^{2}}y^{2}\sqrt{1-|y|^{2}}\\ -\frac{y^{1}y^{2}}{a^{2}}&1-(y^{2})^{2}-\frac{a^{2}-1}{a^{2}}(y^{1})^{2}&-y^{2}y^{3}-\frac{a^{2}-1}{a^{2}}y^{1}\sqrt{1-|y|^{2}}\\ -y^{1}y^{3}+\frac{a^{2}-1}{a^{2}}y^{2}\sqrt{1-|y|^{2}}&-y^{2}y^{3}-\frac{a^{2}-1}{a^{2}}y^{1}\sqrt{1-|y|^{2}}&\frac{a^{2}-1}{a^{2}}(y^{2})^{2}+\frac{a^{2}-1}{a^{2}}(y^{2})^{2}-\frac{1-(y^{3})^{2}}{a^{2}}\\ \end{pmatrix}$. (C.5)
@@
    $\widetilde{V}_{1}:=V_{1}\,,\qquad\widetilde{V}_{2}:=V_{2}\,,\qquad\widetilde{V}_{3}:=\frac{1}{a}V_{3}$ (C.6)
    are orthonormal with respect to the metric $g$ defined by (C.5).
@@
    The operator $B$ defined in accordance with (C.2) is the Laplace–Beltrami operator (henceforth denoted by $\Delta$) on the Berger sphere.
@@
    $\operatorname{curl}_{s}=\begin{pmatrix}\frac{2}{a}&-\frac{1}{a}V_{3}&V_{2}\\ \frac{1}{a}V_{3}&\frac{2}{a}&-V_{1}\\ -V_{2}&V_{1}&2a\\ \end{pmatrix}$. (C.7)
    Here the subscript s\,s\, indicates that the operator appearing in (C.7) acts on 3-columns of scalar functions. The latter can be obtained by projecting 1-forms onto the orthonormal framing (C.6). The operators $\operatorname{curl}$ and $\operatorname{curl}_{s}$ are related as
@@
    See also the projection operator $\mathbf{S}$ in [16, Section 5.1].
C.2 The spectrum
Theorem C.2.
The eigenvalues of the operator $\Delta$ on the Berger 3-sphere are
$n(n+2)+\left(a^{-2}-1\right)\left(n-2l\right)^{2}\,,\quad n=0,1,2,3,\dots,\quad 0\leq l\leq\left\lfloor\frac{n}{2}\right\rfloor$ (C.8)